The excess of electron that must be placed on each of two small spheres spaced 3 cm apart if the force of repulsion between the spheres is to be 10^-19 N is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

let the charge on spheres is q

so F= 1/4∏ε qq/.032

  given F= 10-19

       1/4∏ε qq/.032 =10-19

 9 *109 * q2 /.032 =10-19 

               q=10-16

1 electron has  1.6 * 10-19 C  

 ne =q

 n=q/e =10-16/1.6 * 10-19 C

           =625

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin